Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 36     Entry time: 12 Mar 2025, 23:14
Author: Raphael de Wijn 
Group: SPB 
Subject: Shift summary 

SPB/SFX shift summary user experiment 6933 (setup day)

X-ray delivery
-9.6-9.72 keV with Hirex monitoring, 1.4-1.8 mJ/pulse, 352 pulses/train at 1.1 MHz pulse repetition rate, short pulses

Optical laser delivery
-ns laser for illumination

Achievements/Observations

  • Beamline aligned
  • Energy scan with Zn done and energy set around 9.72 keV (120 eV offset compared to the accelerator value)
  • Aligned and could see scattering from silver cubes
ELOG V3.1.4-7c3fd00